three times the measure of the supplement of an angle is equal to eight times the measure of its complement. find the angle, its complement, and its supplement.
Answer by BurningFlame_marc(77)

Hello!
First,let x be the angle,180-x be the supplement,9-x be the complement.then,we need to do the equation.
Solution:
3(180-x)=8(90-x)
540-3x=720-8x
540=720-5x
-180=-5x
36=x
x=36
Final Answer:
the angle is 36,supplement is 144,complement is 54.
Thank You!!!